# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 4.5.1. The Limit Measure
(4.12) $\nu_{l}(\mathcal{S}_{1/(C_{1}\beta_{l})}(B(\mathbf{0},\Lambda/6))\setminus\bigl([-1,1]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R}\bigl))\lesssim\frac{1}{l}.$
We take the weak limit (via a subsequence) to a non-degenerate measure $\nu$ supported in $1,1]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R}$.
The cylindrically vampiric property of $\nu$: We next examine how the square function coefficients behave under the stretch mapping. For all $l$ sufficiently large and all $\psi\in\mathcal{F}_{l}$, owing to (4.7) we have
$\int_{B(\mathbf{0},\Lambda)}\bigg|\int_{\mathbb{R}^{n+1}}(X-Y)\psi(\mathbf{Y}^{-1}\cdot\mathbf{X})d\mu_{l}(\mathbf{Y})\bigg|^{2}d\mu_{l}(\mathbf{X})\lesssim\frac{\beta_{l}^{2}}{l}.$
@@
$\int_{\mathcal{R}^{l}(B(\mathbf{0},\Lambda)\cap L_{l,C_{1}\beta_{l}})}\bigg|\int_{\mathbb{R}^{n+1}}\frac{(X^{\prime}-Y^{\prime})}{\beta_{l}}\cdot(\psi\circ(\mathcal{O}^{l})^{-1})(X^{\prime}-Y^{\prime},X^{\prime\prime}-Y^{\prime\prime},t-s+\eta((X^{\prime},X^{\prime\prime}),(Y^{\prime},Y^{\prime\prime})))\\ d\mu_{l}((\mathcal{R}^{l})^{-1}(Y^{\prime},Y^{\prime\prime},s))\bigg|^{2}d\mu_{l}((\mathcal{R}^{l})^{-1}(X^{\prime},X^{\prime\prime},t))\lesssim\frac{1}{l}.$
@@
$\int_{B(0,\Lambda/2)}\bigg|\int_{\mathbb{R}^{n+1}}(X^{\prime}-Y^{\prime})\cdot(\psi\circ(\mathcal{O}^{l})^{-1})\big(C_{1}\beta_{l}(X^{\prime}-Y^{\prime}),X^{\prime\prime}-Y^{\prime\prime},t-s+\\ \eta((C_{1}\beta_{l}X^{\prime},X^{\prime\prime}),(C_{1}\beta_{l}Y^{\prime},Y^{\prime\prime}))\big)d\nu_{l}(Y^{\prime},Y^{\prime\prime},s)\bigg|^{2}d\nu_{l}(X^{\prime},X^{\prime\prime},t)\lesssim\frac{1}{l}$
Hence, the measure $\nu$ satisfies (using that $\eta((0,X^{\prime\prime}),(0,Y^{\prime\prime}))=0$ and the continuity of $\eta$)
(4.13) $\int_{\mathbb{R}^{n+1}}(X^{\prime}-Y^{\prime})(\psi\circ\mathcal{O}^{-1})(0,X^{\prime\prime}-Y^{\prime\prime},t-s)d\nu(Y^{\prime},Y^{\prime\prime},s)=0,$
(4.14)  $\;\;\text{ for every }(X^{\prime},X^{\prime\prime},t)\in\operatorname{supp}(\nu)\cap B(\mathbf{0},\Lambda/2)\text{ and }\psi\in\operatorname{Lip}_{0}^{\operatorname{even}}(B(\mathbf{0},\Lambda)).$
@@
The support of $\nu$ is a graph: Consider the minimizing measures $\sigma_{l}(\,\cdot\,\mathbf{X}_{l})\in\mathcal{M}_{k}$ for the quantity $\alpha_{k,\mu_{l}}(Q_{l})$, where $\sigma_{l}\in\mathcal{M}_{k}$. Let $\sigma_{l}=\mathcal{H}^{k}|_{P_{l}}\times\gamma_{l}$, where $P_{l}\subset\mathbb{R}^{n}$ is a $k$-plane containing the origin, and $\gamma_{l}$ is a $2$-ADR measure on $\mathbb{R}$ containing 0 (and whose ADR constant may be chosen depending on the ADR constant $C_{0}$ of $\mu_{\ell}$). After passing to a subsequence, we are free to assume
$\sigma_{l}\rightharpoonup\sigma=\mathcal{H}^{k}|_{P}\times\gamma$
for some vertical $k$-plane $P$ containing the origin and some $2$-ADR measure $\gamma$ containing 0. After passing to another subsequence, we can assume that
$\sigma_{l}(\,\cdot\,\mathbf{X}_{l})\rightharpoonup\sigma(\,\cdot\,\mathbf{X})$
@@
Because $\mathbf{0}\in\operatorname{supp}(\mu)$ and $\beta_{\mu}(\mathbf{0},\Lambda)=0$, it is easy to see that $L$ is a vertical $k$-plane containing the origin, and moreover that $\operatorname{supp}(\mu)\cap B_{\Lambda}(\mathbf{0})\subset L$. But since $\alpha_{k,\mu}(\mathbf{0},1)=0$, it must be the case that $\operatorname{supp}(\sigma(\,\cdot\,\mathbf{X}))\subset L$. Write $\mathbf{X}=(X,t)\in\mathbb{R}^{n}\times\mathbb{R}$ and suppose, for contradiction, that $X\notin P$. Then $\operatorname{supp}(\sigma(\,\cdot\,\mathbf{X}))$ cannot be a vertical plane through the origin, as the spatial projection of its support is $P+X$. But this contradicts that $\operatorname{supp}(\sigma(\,\cdot\,\mathbf{X}))\subset L$. So, $X\in P$, and it easily follows that $\operatorname{supp}(\sigma(\,\cdot\,\mathbf{X}))\subset P\times\mathbb{R}$. Hence, $L=P\times\mathbb{R}$. Moreover, using that $\eta(X,Y)=0$ whenever $X$ and $Y$ are in the same $k$-dimensional subspace of $\mathbb{R}^{n}$, we see that
$\sigma(\,\cdot\,\mathbf{X})=\sigma(\,\cdot\,(0,t))=\mathcal{H}^{k}|_{P}\times\gamma(\,\cdot+t).$
In the sequel, we abuse notation and relabel the measure $\gamma(\,\cdot\,+t)$ as $\gamma$.
@@
(4.15) $\int_{B_{\Lambda/6}(\mathbf{0})}f(0,X^{\prime\prime},t)d\big(\nu-(\mathcal{H}^{k}|_{\{0\}\times\mathbb{R}^{k}}\times\gamma)\big)=0.$
@@
$\int_{\mathbb{R}^{n+1}}f(X^{\prime},X^{\prime\prime},t)d\mu_{\ell}\circ(\mathcal{R}^{\ell})^{-1}(X^{\prime},X^{\prime\prime},t)=\int_{\mathbb{R}^{n+1}}f(\beta_{\ell}X^{\prime},X^{\prime\prime},t)d\nu_{\ell}(X^{\prime},X^{\prime\prime},t)$
But now, if $f\in\operatorname{Lip}_{0}(\mathbf{0},\Lambda)$, then
$\lim_{\ell\to\infty}\int_{\mathbb{R}^{n+1}}f(\beta_{\ell}X^{\prime},X^{\prime\prime},t)d\nu_{\ell}(X^{\prime},X^{\prime\prime},t)=\int_{\mathbb{R}^{n+1}}f(0,X^{\prime\prime},t)d\nu(X^{\prime},X^{\prime\prime},t).$
@@
$\Phi(\mathbf{X})=\inf_{\mathbf{Y}=(0,Y^{\prime\prime},s)\in A}\Bigl\{\varphi(Y^{\prime\prime})\xi(s)+L\|\mathbf{Y}^{-1}\cdot\mathbf{X}\|\Bigl\},$
where $L$ is the Lipschitz constant of $\varphi\xi$ on $A$. Since $\varphi$ is even (and $\|(-Y,-Y^{\prime\prime},s))^{-1}\cdot(-X^{\prime},-X^{\prime\prime},t)\|=\|(Y^{\prime},Y^{\prime\prime},s)^{-1}\cdot(X^{\prime},X^{\prime\prime},t)\|$, we have that $\Phi(-X,t)=\Phi(X,t)$. Multiplying by a suitable cut-off function completes the proof. ∎
Observe that $1,1]\times B^{(k)}_{\Lambda/12}(\mathbf{0})\times[-\Lambda^{2}/144,\Lambda^{2}/144]\subset B_{\Lambda/6}(\mathbf{0})$ if $\Lambda$ is large enough, so, since $\nu$ is supported in $1,1]\times\mathbb{R}^{k}\times\mathbb{R}$, we infer from (4.15) and Lemma 4.10, that
(4.16) $\int\limits_{\mathbb{R}^{n-k}\times B^{(k)}_{\Lambda/12}(0)\times[-\Lambda^{2}/144,\Lambda^{2}/144]}\!\!\!\!\!\!\!\!\!\!\varphi(X^{\prime\prime})\xi(t)d[\nu(X^{\prime},X^{\prime\prime},t)-\mathcal{H}^{k}_{|\{0\}\times\mathbb{R}^{k}}(X^{\prime},X^{\prime\prime})d\gamma(t)]=0.$
@@
(4.17) $\nu(\mathbb{R}^{n-k}\times E\times I)=\nu([-1,1]^{n-k}\times E\times I)=\mathcal{H}^{k}(E)\gamma(I).$
The display (4.17) implies that, for each $X^{\prime\prime}\in B^{(k)}_{\Lambda/12}(\mathbf{0})$ and $t\in[-\Lambda^{2}/144,\Lambda^{2}/144]$, there exists (possibly more than one) $X^{\prime}\in[-1,1]^{n-k}$ with $(X^{\prime},X^{\prime\prime},t)\in\operatorname{supp}(\nu)$. Choose $X^{\prime\prime}\in B^{(k)}_{\Lambda/24}(\mathbf{0})$, $t\in[-\Lambda^{2}/288,\Lambda^{2}/288]$, and select $X^{\prime}\in[-1,1]^{n-k}$ with $(X^{\prime},X^{\prime\prime},t)\in\operatorname{supp}(\nu)$.
For $Z^{\prime\prime}\in B_{\Lambda/24}(\mathbf{0})$, the points $X^{\prime\prime}+Z^{\prime\prime}\in B_{\Lambda/12}(\mathbf{0})$ and $X^{\prime\prime}-Z^{\prime\prime}\in B_{\Lambda/12}(\mathbf{0})$. Also select $u\in[-\Lambda^{2}/288,\Lambda^{2}/288]$. For sufficiently small $\varepsilon>0$, Lemma 4.10 ensures that the indicator function of the set $\{0\}\times(\chi_{B_{\varepsilon}^{(k)}(Z^{\prime\prime})}\cup\chi_{B_{\varepsilon}^{(k)}(-Z^{\prime\prime})})\times(u-\varepsilon,u+\varepsilon)$ can be approximated by functions in $\operatorname{Lip}_{0}^{\operatorname{even}}(B_{\Lambda/6}(\mathbf{0}))$. Denote by
$E_{\varepsilon}(Y^{\prime\prime},u)=[-1,1]^{n-k}\times B_{\varepsilon}^{(n-k)}(Y^{\prime\prime})\times(u-\varepsilon,u+\varepsilon).$
Then using (4.13) we obtain
$\int_{E_{\varepsilon}(X^{\prime\prime}+Z^{\prime\prime},t+u)}(X^{\prime}-Y^{\prime})d\nu(Y^{\prime},Y^{\prime\prime},s)+\int_{E_{\varepsilon}(X^{\prime\prime}-Z^{\prime\prime},t+u)}(X^{\prime}-Y^{\prime})d\nu(Y^{\prime},Y^{\prime\prime},s)=0.$
@@
(4.18) $\begin{split}X^{\prime}=\frac{1}{2\mathcal{H}^{(k)}(B(0,\varepsilon))\sigma((s-\varepsilon,s+\varepsilon))}\bigg(&\int_{E_{\varepsilon}(X^{\prime\prime}+Z^{\prime\prime},t+u)}Y^{\prime}d\nu(Y^{\prime},Y^{\prime\prime},s)\\ &+\int_{E_{\varepsilon}(X^{\prime\prime}-Z^{\prime\prime},t+u)}Y^{\prime}d\nu(Y^{\prime},Y^{\prime\prime},s)\bigg).\end{split}$
This formula determines $X^{\prime}$ uniquely, and hence $\operatorname{supp}(\nu)\cap([-1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]$ given by $\{(A(X^{\prime\prime},t),X^{\prime\prime},t)\}$. But then we also have (letting $\varepsilon$ tend to zero)
(4.19) $A(X^{\prime\prime},t)=\frac{1}{2}\bigl(A(X^{\prime\prime}+Z^{\prime\prime},t+u)+A(X^{\prime\prime}-Z^{\prime\prime},t+u)\bigl)$
for every $X^{\prime\prime},Z^{\prime\prime}\in B^{(k)}_{\Lambda/24}(\mathbf{0})$ and $t,u\in[-\Lambda^{2}/288,\Lambda^{2}/288]$. Hence, considering this identity with $u=0$ ensures that $A(\,\cdot\,,t)$ is an affine function. But then, (4.19) now ensures that $A(\,\cdot\,,t)=A(\,\cdot\,,t+u)$ for all $u\in[-\Lambda^{2}/288,\Lambda^{2}/288]$. Therefore, there is an affine function $A:B^{(k)}_{\Lambda^{2}/288}(0)\to[-1,1]^{n-k}$ such that $\operatorname{supp}(\nu)\cap[-1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]$ equals
$\{(A(X^{\prime\prime}),X^{\prime\prime},t):(X^{\prime\prime},t)\in B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]\}$
@@
$1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]\supset B(\mathbf{0},3).$
@@
$\int_{B(\mathbf{0},3)}|X^{\prime}-A(X^{\prime\prime})|^{2}d\nu(X,t)=0,$
@@
$\int_{B(\mathbf{0},3)\cap\{|X^{\prime}|\leq C_{1}\beta_{\ell}\}}\frac{|X^{\prime}-\beta_{\ell}AX^{\prime\prime}|^{2}}{\beta_{k}^{2}}d\mu_{\ell}\circ(\mathcal{R}^{(\ell)})^{-1}(X^{\prime},X^{\prime\prime},t)\to 0.$
If we define $\widetilde{L}_{\ell}=(\mathcal{R}^{\ell})^{-1}\{\beta_{\ell}A(X^{\prime\prime}),X^{\prime\prime},t)\}$, then as $L_{\ell}$ passes close to the origin
$\int_{B(\mathbf{0},2)\cap L_{\ell,C_{1}\beta_{\ell}}}\frac{\operatorname{dist}(\mathbf{X},\widetilde{L}_{\ell})^{2}}{\beta_{k}^{2}}d\mu_{\ell}(\mathbf{X})\to 0.$
But also, $\widetilde{L}_{\ell}\cap B(\mathbf{0},2)$ lies in the $C_{1}\beta_{\ell}$-neighbourhood of $L_{\ell}\cap B(\mathbf{0},2)$, so (4.9) ensures that
$\int_{B(\mathbf{0},2)\setminus L_{l,C_{0}\beta_{l}}}\frac{\text{dist}(\mathbf{X},\widetilde{L}_{\ell})^{2}}{\beta_{l}^{2}}d\mu_{l}\xrightarrow{l\rightarrow\infty}0.$

### v2

@@ 4.5.1. The Limit Measure
(4.14) $\sigma_{\ell}(\mathcal{S}_{1/(C_{1}\beta_{\ell})}(B_{\Lambda/6}(\mathbf{0}))\setminus\bigl([-1,1]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R}\bigl))\lesssim\frac{1}{\ell}.$
We take the weak limit (via a subsequence) to a non-degenerate measure $\sigma$ supported in $1,1]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R}$.
The cylindrically vampiric property of $\sigma$: We next examine how the square function coefficients behave under the stretch mapping. For all $\ell$ sufficiently large and all $\psi\in\mathcal{F}_{\ell}$, owing to (4.9) we have
$\int_{B_{\Lambda}(\mathbf{0})}\bigg|\int_{\mathbb{R}^{n+1}}(X-Y)\psi(\mathbf{Y}^{-1}\cdot\mathbf{X})d\mu_{\ell}(\mathbf{Y})\bigg|^{2}d\mu_{\ell}(\mathbf{X})\lesssim\frac{\beta_{\ell}^{2}}{\ell}.$
@@
$\int_{\mathcal{R}^{\ell}(B_{\Lambda}(\mathbf{0})\cap L_{\ell,C_{1}\beta_{\ell}})}\bigg|\int_{\mathbb{R}^{n+1}}\frac{(X^{\prime}-Y^{\prime})}{\beta_{\ell}}\cdot(\psi\circ(\mathcal{O}^{\ell})^{-1})(X^{\prime}-Y^{\prime},X^{\prime\prime}-Y^{\prime\prime},t-s+\eta((X^{\prime},X^{\prime\prime}),(Y^{\prime},Y^{\prime\prime})))\\ d\mu_{\ell}((\mathcal{R}^{\ell})^{-1}(Y^{\prime},Y^{\prime\prime},s))\bigg|^{2}d\mu_{\ell}((\mathcal{R}^{\ell})^{-1}(X^{\prime},X^{\prime\prime},t))\lesssim\frac{1}{\ell}.$
@@
$\int_{B_{\Lambda/2}(\mathbf{0})}\bigg|\int_{\mathbb{R}^{n+1}}(X^{\prime}-Y^{\prime})\cdot(\psi\circ(\mathcal{O}^{\ell})^{-1})\big(C_{1}\beta_{\ell}(X^{\prime}-Y^{\prime}),X^{\prime\prime}-Y^{\prime\prime},t-s+\\ \eta((C_{1}\beta_{\ell}X^{\prime},X^{\prime\prime}),(C_{1}\beta_{\ell}Y^{\prime},Y^{\prime\prime}))\big)d\sigma_{\ell}(Y^{\prime},Y^{\prime\prime},s)\bigg|^{2}d\sigma_{\ell}(X^{\prime},X^{\prime\prime},t)\lesssim\frac{1}{\ell}.$
Hence, the measure $\sigma$ satisfies (using that $\eta((0,X^{\prime\prime}),(0,Y^{\prime\prime}))=0$ and the continuity of $\eta$)
(4.15) $\int_{\mathbb{R}^{n+1}}(X^{\prime}-Y^{\prime})(\psi\circ\mathcal{O}^{-1})(0,X^{\prime\prime}-Y^{\prime\prime},t-s)d\sigma(Y^{\prime},Y^{\prime\prime},s)=0,$
(4.16)  $\;\;\text{ for every }(X^{\prime},X^{\prime\prime},t)\in\operatorname{supp}(\sigma)\cap B_{\Lambda/2}(\mathbf{0})\text{ and }\psi\in\operatorname{Lip}_{0}^{\operatorname{even}}(B_{\Lambda}(\mathbf{0})).$
@@
The support of $\sigma$ is a graph: Consider the minimizing measures $\zeta_{\ell}(\,\cdot\,\mathbf{X}_{\ell})\in\mathcal{M}_{k}$ for the quantity $\alpha_{k,\mu_{\ell}}(Q_{\ell})$, where $\zeta_{\ell}\in\mathcal{M}_{k}$. Let $\zeta_{\ell}=\mathcal{H}^{k}|_{P_{\ell}}\times\gamma_{\ell}$, where $P_{\ell}\subset\mathbb{R}^{n}$ is a $k$-plane containing the origin, and $\gamma_{\ell}$ is a $2$-ADR measure on $\mathbb{R}$ containing 0 (and whose ADR constant may be chosen depending on the ADR constant of $\mu_{\ell}$, which is uniformly bounded over $\ell$). After passing to a subsequence, we are free to assume
$\zeta_{\ell}\rightharpoonup\zeta=\mathcal{H}^{k}|_{P}\times\gamma$
for some vertical $k$-plane $P$ containing the origin and some $2$-ADR measure $\zeta$ containing 0. After passing to another subsequence, we can assume that
$\zeta_{\ell}(\,\cdot\,\mathbf{X}_{\ell})\rightharpoonup\zeta(\,\cdot\,\mathbf{X})$
@@
Because $\mathbf{0}\in\operatorname{supp}(\mu)$ and $\beta_{\mu}(\mathbf{0},\Lambda)=0$, it is easy to see that $L$ is a vertical $k$-plane containing the origin, and moreover that $\operatorname{supp}(\mu)\cap B_{\Lambda}(\mathbf{0})\subset L$. Since $\alpha_{k,\mu}(\mathbf{0},\Lambda)=0$, it must be the case that $\operatorname{supp}(\zeta(\,\cdot\,\mathbf{X}))\subset L$ as well. Write $\mathbf{X}=(X,t)\in\mathbb{R}^{n}\times\mathbb{R}$ and suppose, for contradiction, that $X\notin P$. Then $\operatorname{supp}(\zeta(\,\cdot\,\mathbf{X}))$ cannot be a vertical plane through the origin, as the spatial projection of its support is PXP-X. But this contradicts that $\operatorname{supp}(\zeta(\,\cdot\,\mathbf{X}))\subset L$. So, $X\in P$, and it easily follows that $\operatorname{supp}(\zeta(\,\cdot\,\mathbf{X}))\subset P\times\mathbb{R}$. Hence, $L=P\times\mathbb{R}$. Moreover, using that $\eta(X,Y)=0$ whenever $X$ and $Y$ are in the same $k$-dimensional subspace of $\mathbb{R}^{n}$, we see that
$\zeta(\,\cdot\,\mathbf{X})=\zeta(\,\cdot\,(0,t))=\mathcal{H}^{k}|_{P}\times\gamma(\,\cdot+t).$
In the sequel, we abuse notation and relabel the measure $\gamma(\,\cdot\,+t)$ as $\gamma$, as we will use $t$ to denote other points of $\mathbb{R}$.
@@
(4.17) $\int_{B_{\Lambda/6}(\mathbf{0})}f(0,X^{\prime\prime},t)d\big(\sigma-(\mathcal{H}^{k}|_{\{0_{n-k}\}\times\mathbb{R}^{k}}\times\gamma)\big)=0.$
@@
$\int_{\mathbb{R}^{n+1}}f(X^{\prime},X^{\prime\prime},t)d\mu_{\ell}\circ(\mathcal{R}^{\ell})^{-1}(X^{\prime},X^{\prime\prime},t)=\int_{\mathbb{R}^{n+1}}f(C_{1}\beta_{\ell}X^{\prime},X^{\prime\prime},t)d\sigma_{\ell}(X^{\prime},X^{\prime\prime},t)$
But now, if $f\in\operatorname{Lip}_{0}(B_{\Lambda}(\mathbf{0}))$, then
$\lim_{\ell\to\infty}\int_{\mathbb{R}^{n+1}}f(C_{1}\beta_{\ell}X^{\prime},X^{\prime\prime},t)d\sigma_{\ell}(X^{\prime},X^{\prime\prime},t)=\int_{\mathbb{R}^{n+1}}f(0,X^{\prime\prime},t)d\sigma(X^{\prime},X^{\prime\prime},t).$
@@
$\Phi(\mathbf{X})=\inf_{\mathbf{Y}=(0,Y^{\prime\prime},s)\in A}\Bigl\{\varphi(Y^{\prime\prime})\xi(s)+\lambda\|\mathbf{Y}^{-1}\cdot\mathbf{X}\|\Bigl\},$
where $\lambda$ is the Lipschitz constant of $\varphi\xi$ on $A$. Since $\varphi$ is even (and $\|(-Y,-Y^{\prime\prime},s))^{-1}\cdot(-X^{\prime},-X^{\prime\prime},t)\|=\|(Y^{\prime},Y^{\prime\prime},s)^{-1}\cdot(X^{\prime},X^{\prime\prime},t)\|$, we have that $\Phi(-X,t)=\Phi(X,t)$. Multiplying by a suitable cut-off function completes the proof. ∎
Observe that $1,1]^{n-k}\times B^{(k)}_{\Lambda/12}(0)\times[-\Lambda^{2}/144,\Lambda^{2}/144]\subset B_{\Lambda/6}(\mathbf{0})$ for $\Lambda\geq 3000$, so, since $\sigma$ is supported in $1,1]\times\mathbb{R}^{k}\times\mathbb{R}$, we infer from (4.17) and Lemma 4.10, that for $\varphi$, $\xi$ as in Lemma 4.10
(4.18) $\int\limits_{\mathbb{R}^{n-k}\times B^{(k)}_{\Lambda/12}(0)\times[-\Lambda^{2}/144,\Lambda^{2}/144]}\!\!\!\!\!\!\!\!\!\!\varphi(X^{\prime\prime})\xi(t)d[\sigma(X^{\prime},X^{\prime\prime},t)-\mathcal{H}^{k}_{|\{0\}\times\mathbb{R}^{k}}(X^{\prime},X^{\prime\prime})\gamma(t)]=0.$
@@
(4.19) $\sigma(\mathbb{R}^{n-k}\times E\times I)=\sigma([-1,1]^{n-k}\times E\times I)=\mathcal{H}^{k}(E)\gamma(I).$
Thus, for each $X^{\prime\prime}\in B^{(k)}_{\Lambda/12}(\mathbf{0})$ and $t\in[-\Lambda^{2}/144,\Lambda^{2}/144]\cap\operatorname{supp}(\gamma)$, there exists (at least one) $X^{\prime}\in[-1,1]^{n-k}$ such that $(X^{\prime},X^{\prime\prime},t)\in\operatorname{supp}(\sigma).$
Choose $(X^{\prime},X^{\prime\prime},t)\in([-1,1]^{n-k}\times B_{\Lambda/24}^{(k)}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288])\cap\operatorname{supp}(\sigma)$ (in particular, using 4.19, $t\in\operatorname{supp}(\gamma)$). Denote by
$E_{\varepsilon}(\zeta^{\prime\prime},s)=[-2,2]^{n-k}\times B_{\varepsilon}^{(k)}(\zeta^{\prime\prime})\times(s-\epsilon,s+\epsilon).$
Choose $u\in\operatorname{supp}(\gamma)\cap[-\Lambda^{2}/288,\Lambda^{2}/288]$. Then, with $Z^{\prime\prime}\in B^{(k)}_{\Lambda/24}(0)$, applying (4.15) with (Lipschitz approximations of) the kernel
$\chi_{{}_{E_{\varepsilon}(Z^{\prime\prime},t+u)}}+\chi_{{}_{E_{\varepsilon}(-Z^{\prime\prime},t+u)}},$
we obtain, for $\epsilon$ sufficiently small
$\int\big(X^{\prime}-Y^{\prime})\big(\chi_{{}_{E_{\varepsilon}(Z^{\prime\prime},t+u)}}(X^{\prime}-Y^{\prime},X^{\prime\prime}-Y^{\prime\prime},t-s)\\ +\chi_{{}_{E_{\varepsilon}(-Z^{\prime\prime},t+u)}}(X^{\prime}-Y^{\prime},X^{\prime\prime}-Y^{\prime\prime},t-s)\big)d\sigma(Y^{\prime\prime},Y^{\prime},s)\\ =\int_{E_{\varepsilon}(X^{\prime\prime}-Z^{\prime\prime},u)}(X^{\prime}-Y^{\prime})d\sigma(Y^{\prime},Y^{\prime\prime},s)+\int_{E_{\varepsilon}(X^{\prime\prime}+Z^{\prime\prime},u)}(X^{\prime}-Y^{\prime})d\sigma(Y^{\prime},Y^{\prime\prime},s)=0.$
@@
(4.20) $\begin{split}X^{\prime}=\frac{1}{2\mathcal{H}^{k}(B^{(k)}_{\varepsilon}(0))\gamma((u-\varepsilon,u+\varepsilon))}\bigg(&\int_{E_{\varepsilon}(X^{\prime\prime}+Z^{\prime\prime},u)}Y^{\prime}d\sigma(Y^{\prime},Y^{\prime\prime},s)\\ &+\int_{E_{\varepsilon}(X^{\prime\prime}-Z^{\prime\prime},u)}Y^{\prime}d\sigma(Y^{\prime},Y^{\prime\prime},s)\bigg).\end{split}$
Notice that if $\tilde{X^{\prime}}$ is another point such that $(\tilde{X^{\prime}},X^{\prime\prime},t)\in\mathcal{S}\cap\operatorname{supp}(\sigma)$, then $\tilde{X^{\prime}}$ satisfies the same inequality above, and hence $X^{\prime}=\tilde{X^{\prime}}$. Hence, $\operatorname{supp}(\sigma)\cap([-1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288])$ is a subset of a graph $\{(A(X^{\prime\prime},t),X^{\prime\prime},t)\}_{(X^{\prime\prime},t)\in\mathbb{R}^{k}\times\mathbb{R}}$. But then we also have (letting $\varepsilon$ tend to zero)
(4.21) $A(X^{\prime\prime},t)=\frac{1}{2}\bigl(A(X^{\prime\prime}+Z^{\prime\prime},u)+A(X^{\prime\prime}-Z^{\prime\prime},u)\bigl)$
for every $X^{\prime\prime},Z^{\prime\prime}\in B^{(k)}_{\Lambda/24}(\mathbf{0})$ and $t,u\in[-\Lambda^{2}/288,\Lambda^{2}/288]\cap\operatorname{supp}(\gamma)$. 4.21 tells us two things. First, taking $t=u$, we see that for every $t\in\operatorname{supp}(\gamma)\cap[-\Lambda^{2}/288,\Lambda^{2}/288]$, $A(\cdot,t)$ is an affine functions in the spatial components. But, we also will have that the $A(X^{\prime\prime},t)=A(X^{\prime\prime},u)$ for all $t,u\in\operatorname{supp}(\gamma)\cap[-\Lambda^{2}/288,\Lambda^{2}/288]$, so we can take $A$ to be an affine functions which is independent of time. In other words, there is an affine function independent of time $A:B^{(k)}_{\Lambda^{2}/288}(0)\times\mathbb{R}\to[-1,1]^{n-k}$ such that $\operatorname{supp}(\sigma)\cap\big([-1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]\big)$ equals
$\big\{(A(X^{\prime\prime}),X^{\prime\prime},t):(X^{\prime\prime},t)\in B^{(k)}_{\Lambda/24}(0)\times\big([-\Lambda^{2}/288,\Lambda^{2}/288])\cap\operatorname{supp}(\gamma)\big)\big\}$
@@
$1,1]^{n-k}\times B^{(k)}_{\Lambda/24}(0)\times[-\Lambda^{2}/288,\Lambda^{2}/288]\supset B_{1}(\mathbf{0}).$
@@
$\int_{B_{1}(\mathbf{0})}|X^{\prime}-A(X^{\prime\prime},t)|^{2}d\sigma(X,t)=0,$
@@
$\int_{B_{1}(\mathbf{0})\cap\{|X^{\prime}|\leq C_{1}\beta_{\ell}\}}\frac{|X^{\prime}-\beta_{\ell}A(X^{\prime\prime},t)|^{2}}{\beta_{\ell}^{2}}d\mu_{\ell}\circ(\mathcal{R}^{(\ell)})^{-1}(X^{\prime},X^{\prime\prime},t)\to 0.$
If we define $\widetilde{L}_{\ell}=(\mathcal{R}^{\ell})^{-1}\{(\beta_{\ell}A(X^{\prime\prime},t),X^{\prime\prime},t)\}$, then as $L_{\ell}$ passes close to the origin
$\int_{B_{1}(\mathbf{0})\cap L_{\ell,C_{1}\beta_{\ell}}}\frac{\operatorname{dist}(\mathbf{X},\widetilde{L}_{\ell})^{2}}{\beta_{\ell}^{2}}d\mu_{\ell}(\mathbf{X})\to 0.$
The display above implies that, for large $\ell$, $\widetilde{L}_{\ell}\cap B_{2}(\mathbf{0})$ lies in the $C_{1}\beta_{\ell}$-neighbourhood of $L_{\ell}\cap B_{2}(\mathbf{0})$. Hence, (4.11) ensures that
$\int_{B_{1}(\mathbf{0})\setminus L_{\ell,C_{1}\beta_{\ell}}}\frac{\text{dist}(\mathbf{X},\widetilde{L}_{\ell})^{2}}{\beta_{\ell}^{2}}d\mu_{\ell}\xrightarrow{\ell\rightarrow\infty}0.$